OpenStudy (anonymous):

Solve. x – 9 = –15 24 6 –6 –24

OpenStudy (anonymous):

i figured t would be D

OpenStudy (anonymous):

add 9 on both sides then you get 6

OpenStudy (anonymous):

-6 *

OpenStudy (anonymous):

because its subtracting to u have to flip it nd change it 2 24

OpenStudy (haseeb96):

x-9=-15 x=-15+9 x=-6

OpenStudy (anonymous):

@Starr_DynastyT you are incorrect